Banyan Co.’s common stock currently sells for $53.75 per share. The growth rate is a constant 13.5%, and the company has an expected dividend yield of 3%. The expected long-run dividend payout ratio is 25%, and the expected return on equity (ROE) is 18%. New stock can be sold to the public at the current price, but a flotation cost of 15% would be incurred. What would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations.

%
Answer #1

Current Share Price = $53.75

Expected Dividend Yield = D1/Stock Price

Expected Dividend Yield = 3%

So,

D1 = 0.03(53.75) = $1.6125

So,

As per Constant Growth Model,

Stock Price(1 - f) = D1/(r - g)

r = 1.6125/53.75(1 - 0.15) + 0.1350

r = 17.03%

So,

Cost of New Equity = 17.03%